# SBP-FDEC: Summation-by-Parts Finite Difference Exterior Calculus

### v2

@@ 2.2.3 Divergence Operator
$\begin{split}f(x_{i},y_{j},z_{z})=&\sum_{s=1}^{N}\sum_{t=1}^{N}\sum_{r=1}^{N}(\tilde{w}_{s,t,r}^{x}-\tilde{w}_{s-1,t,r}^{x}+\tilde{w}_{s,t,r}^{y}\\ &-\tilde{w}_{s,t-1,r}^{y}+\tilde{w}_{s,t,r}^{z}-\tilde{w}_{s,t,r-1}^{z})V_{i,s}V_{j,t}V_{k,r}\\ =&\sum_{s=1}^{N}\sum_{t=1}^{N}\sum_{r=1}^{N}\tilde{w}_{s,t,r}^{x}D_{i,s}V_{j,t}V_{k,r}\\ &+\tilde{w}_{s,t,r}^{y}V_{i,s}D_{j,t}V_{k,r}+\tilde{w}_{s,t,r}^{z}V_{i,s}V_{j,t}D_{k,r}.\end{split}$ (33)
@@
$\textbf{v}^{x}$ $=\begin{bmatrix}v^{x}_{01}\\ \vdots\\ v^{x}_{N1}\\ v^{x}_{02}\\ \vdots\\ v^{x}_{N2}\\ \vdots\\ v^{x}_{NN}\end{bmatrix}\in\mathbb{R}^{N(N+1)},\qquad\textbf{v}^{y}=\begin{bmatrix}v^{y}_{10}\\ \vdots\\ v^{y}_{N0}\\ v^{y}_{11}\\ \vdots\\ v^{y}_{N1}\\ \vdots\\ v^{y}_{NN}\end{bmatrix}\in\mathbb{R}^{N(N+1)},$ (36)
u $=\begin{bmatrix}u_{00}\\ \vdots\\ u_{N0}\\ u_{01}\\ \vdots\\ u_{N1}\\ \vdots\\ u_{NN}\end{bmatrix}\in\mathbb{R}^{(N+1)^{2}},$ (37)
where v is a function for which we require to compute its gradient. Analogously to $\mathbf{u}$, we can define the discrete storage of a function w, for which we need to compute its curl. Using the Kronecker product, we can define the 2D partial derivative operators in matrix form:
$\underline{\Delta}^{2D}_{x}$ $\coloneqq\underline{I}_{N+1}\otimes\underline{\Delta}\in\mathbb{R}^{N(N+1)\times(N+1)^{2}},$ (38)
$\underline{\Delta}^{2D}_{y}$ $\coloneqq\underline{\Delta}\otimes\underline{I}_{N+1}\in\mathbb{R}^{N(N+1)\times(N+1)^{2}},$ (39)
where $\underline{I}_{N+1}$ is the identity matrix of the corresponding dimension.
@@
$\underline{\textbf{grad}}^{x}\left(\textbf{u}\right)$ $=\underline{\Delta}^{2D}_{x}\textbf{u},$ (40)
$\underline{\textbf{grad}}^{y}\left(\textbf{u}\right)$ $=\underline{\Delta}^{2D}_{y}\textbf{u},$ (41)
$\underline{\textbf{curl}}^{x}\left(\textbf{u}\right)$ $=\underline{\Delta}^{2D}_{y}\textbf{u},$ (42)
$\underline{\textbf{curl}}^{y}\left(\textbf{u}\right)$ $=-\underline{\Delta}^{2D}_{x}\textbf{u},$ (43)
$\underline{\text{curl}}\left(\textbf{w}^{x},\textbf{w}^{y}\right)$ $=\underline{\Delta}^{2D}_{x}\textbf{w}^{y}-\underline{\Delta}^{2D}_{y}\textbf{w}^{x},$ (44)
$\underline{\text{div}}\left(\textbf{v}^{x},\textbf{v}^{y}\right)$ $=\underline{\Delta}^{2D}_{x}\textbf{v}^{x}+\underline{\Delta}^{2D}_{y}\textbf{v}^{y},$ (45)
@@
$\underline{\text{curl}}\circ\underline{\textbf{grad}}\left(\textbf{u}\right)$ $=\left(\underline{\Delta}^{2D}_{y}\underline{\Delta}^{2D}_{x}-\underline{\Delta}^{2D}_{x}\underline{\Delta}^{2D}_{y}\right)\textbf{u}=\underline{0},$ (46)
$\underline{\text{div}}\circ\underline{\textbf{curl}}\left(\textbf{u}\right)$ $=\left(\underline{\Delta}^{2D}_{x}\underline{\Delta}^{2D}_{y}-\underline{\Delta}^{2D}_{y}\underline{\Delta}^{2D}_{x}\right)\textbf{u}=\underline{0},$ (47)
@@
With the matrix representation of our derivative operators, we have the commuting 2D dual de Rham complex of Fig. 5, using reduction operators $R^{0},\textbf{R}^{1}$ and $R^{2}$. These operators are defined as the mappings sending functions in the discrete spaces to the vector of their degrees of freedom. We can again evaluate the operators nodally using the histopolation Vandermonde matrix $\underline{V}$ analogously to the 3D case in equations (24) to (33).
@@
Figure 5: Discrete dual de Rham complex in 2D
@@ 2.3 Summation-By-Parts Finite Difference Operators
In the previous section, we assumed the existence of Lagrange functions that fulfill (2) and (3) at grid points $\{x_{i}\}_{i=0}^{N}$ (including the boundary nodes), but we did not discuss how to construct them. In [gerritsma_preprint], Gerritsma uses the Legendre–Gauss–Lobatto points and the Lagrange interpolating polynomials for those points. The corresponding derivative matrix $\underline{D}$ together with the mass matrix $\underline{M}$, which is a diagonal matrix with the Legendre–Gauss–Lobatto quadrature weights on the diagonal, has the summation-by-parts (SBP) property,
@@
As we have pointed out above, we only need to evaluate our basis functions and their derivatives at the grid nodes to construct the mimetic operators defined in the previous section. If we use a collocation quadrature rule when applying Finite Element Exterior Calculus [feec1, feec2] to derive our scheme, the global shape of the basis functions $l_{i}$ does not matter at all, only their nodal function and derivative values are relevant. The central idea of this paper is therefore to not construct basis functions $l_{i}$ at all, but use existing diagonal norm finite difference SBP operators, for example from Strand [strand_sbp], to supply a quadrature rule and nodal derivative data needed for the numerical scheme. We do this explicitly by prescribing the nodal derivatives of the Lagrange functions, that are not known globally, through the derivative matrix according to (13) and using diagonal entries of the norm matrix $\underline{M}$ as weights for a collocation quadrature rule on the grid points.
One of the standard SBP finite difference operators from [strand_sbp], which we will use in this study, is the operator with interior order $4$ and boundary order $2$ with $N+1$ points:
@@
which is defined on an equidistant grid, where $\Delta x$ is the distance between two neighboring grid nodes. In the derivative matrix, the upper right boundary block is repeated in the bottom left with reversed signs and inverted order.
@@ 3.1 Maxwell’s Equations in 2D
Our goal is to derive a mimetic scheme using finite element exterior calculus to preserve the divergence involution (55). For simplicity we will consider a periodic and rectangular domain, using a Cartesian mesh for the discretization. To apply the framework of Finite Element Exterior Calculus, we need the de Rham complexes containing the derivative operators of the PDEs.
@@ 3.2 De Rham Complexes
The Maxwell’s equations have an inherent geometric structure described by the de Rham cochain complex of Fig. 2 and its dual in Fig. 3 in 2D. We define the spaces needed for the complexes by
@@
The complexes are dual in the sense that
@@
$\left<\text{div}\left(\boldsymbol{\Psi}\right),\phi\right$> $=\left<\boldsymbol{\Psi},-\textbf{grad}\left(\phi\right),\right$> $\qquad\forall\boldsymbol{\Psi}\in H\left(\text{div}\right),\phi\in H^{1},$ (59)
where the brackets denote the $L^{2}$ inner product. These equations only hold in the periodic case. In the non-periodic case the boundary integral does not vanish by default, which means we would have to restrict the spaces of one of the two complexes to restore duality.
In Maxwell’s equation we use operators from both complexes, so we have to choose in which spaces to look for E and $B^{z}$. If we choose $\textbf{E}\in H(\text{div})$ and $B^{z}\in H^{1}$, we can solve Ampère’s law and Gauss’s law strongly, while we have to use Stokes’ theorem to solve Faraday’s law weakly, while for $\textbf{E}\in H(\text{curl})$ and $B^{z}\in L^{2}$, we can solve Faraday’s law strongly, but need to solve Ampère’s law and Gauss’s law weakly.
Since the divergence involution for the magnetic field falls away in the 2D transverse electric case and we want to preserve Gauss’ law for the electric field (55), we will choose $\textbf{E}\in H(\text{div})$ and $B^{z}\in H^{1}$, leading to the formulation:
@@
$\left<\frac{\partial{\textbf{E}}}{\partial{t}},\boldsymbol{\Phi}\right>+\left<\textbf{curl}\left(B^{z}\right),\boldsymbol{\Phi}\right$> $=0,\qquad$ $\forall\boldsymbol{\Phi}\in H(\text{div}),$ (61)
@@
The finite element exterior calculus ansatz is to choose a discrete subcomplex of the de Rham complex, meaning we choose $W^{0}\subset H^{1}$, $W^{1}\subset H\left(\text{curl}\right)$, $W^{2}\subset L^{2}$ such that $\textbf{curl}\left(W^{0}\right)\subset W^{1}$ and $\text{div}\left(W^{1}\right)\subset W^{2}$. In addition we need bounded projections $p_{0},\textbf{p}_{1}$ and $p_{2}$, so that the diagram of Fig. 5 commutes.
We then restrict both our ansatz and our test spaces to the corresponding finite dimensional subspaces. This means we again have two ansatz spaces for the electric and magnetic field: We look for a numerical solution $\textbf{E}\in[0,T]\times W^{1}$ and $B^{z}\in[0,T]\times W^{2}$ so that
$\left<\frac{\partial{B^{z}}}{\partial{t}},\psi\right>+\left<\textbf{E},\textbf{curl}\left(\psi\right)\right$> $=0,\qquad$ $\forall\psi\in W^{2},$ (63)
$\left<\frac{\partial{\textbf{E}}}{\partial{t}},\boldsymbol{\Phi}\right>+\left<\textbf{curl}\left(B^{z}\right),\boldsymbol{\Phi}\right$> $=0,\qquad$ $\forall\boldsymbol{\Phi}\in W^{1},$ (64)
$\left<\text{div}\left(\textbf{E}\right),\mu\right$> $=0,\qquad$ $\forall\mu\in W^{0}.$ (65)
Since $\text{div}(\textbf{E})\in W^{0}$ we do get $\text{div}(\textbf{E})\equiv 0$ from (65). We will later show that by solving (63) and (64) with this method, we automatically preserve the divergence involution both in the semi-discrete and fully discrete setting. Therefore we will only consider the evolution equations (63) and (64) when deriving our method.

### v3

@@ 2.2.3 Divergence Operator
$\begin{split}f(x_{i},y_{j},z_{k})=&\sum_{s=1}^{N}\sum_{t=1}^{N}\sum_{r=1}^{N}(\tilde{w}_{s,t,r}^{x}-\tilde{w}_{s-1,t,r}^{x}+\tilde{w}_{s,t,r}^{y}\\ &-\tilde{w}_{s,t-1,r}^{y}+\tilde{w}_{s,t,r}^{z}-\tilde{w}_{s,t,r-1}^{z})V_{i,s}V_{j,t}V_{k,r}\\ =&\sum_{s=0}^{N}\sum_{t=1}^{N}\sum_{r=1}^{N}\tilde{w}_{s,t,r}^{x}D_{i,s}V_{j,t}V_{k,r}\\ +&\sum_{s=1}^{N}\sum_{t=0}^{N}\sum_{r=1}^{N}\tilde{w}_{s,t,r}^{y}V_{i,s}D_{j,t}V_{k,r}\\ +&\sum_{s=1}^{N}\sum_{t=1}^{N}\sum_{r=0}^{N}\tilde{w}_{s,t,r}^{z}V_{i,s}V_{j,t}D_{k,r}.\end{split}$ (34)
@@
[…]
$\underline{u}$ $=\begin{bmatrix}u_{00}\\ \vdots\\ u_{N0}\\ u_{01}\\ \vdots\\ u_{N1}\\ \vdots\\ u_{NN}\end{bmatrix}\in\mathbb{R}^{(N+1)^{2}},$ (38)
where v is a function for which we want to be able to compute the divergence. Analogously to $\mathbf{v}$, we can define the discrete storage of a function w, for which we need to compute its vector-to-scalar curl. Figure 4 shows in which order the DOFs are indexed. Using the Kronecker product, we can define the 2D partial derivative operators in matrix form:
$\underline{\Delta}^{2D}_{x}$ $\coloneqq\underline{I}_{M}\otimes\underline{\Delta}\in\mathbb{R}^{N(N+1)\times M^{2}},$ (39)
$\underline{\Delta}^{2D}_{y}$ $\coloneqq\underline{\Delta}\otimes\underline{I}_{M}\in\mathbb{R}^{N(N+1)\times M^{2}},$ (40)
where $\underline{I}_{M}$ is the $M\times M$ identity matrix. We have $M=N$ if the derivative acts on $U_{l}\otimes U_{l}$ or $M=N+1$ if it acts on $U_{l}\otimes U_{h}$ or $U_{h}\otimes U_{l}$. In an abuse of notation we do note add the value of $M$ as an index of the $\underline{\Delta}^{2D}$ matrices but leave it implied by context
@@
$\underline{\textbf{grad}}^{x}\left(\underline{u}\right)$ $=\underline{\Delta}^{2D}_{x}\underline{u},$ (41)
$\underline{\textbf{grad}}^{y}\left(\underline{u}\right)$ $=\underline{\Delta}^{2D}_{y}\underline{u},$ (42)
$\underline{\textbf{curl}}^{x}\left(\underline{u}\right)$ $=\underline{\Delta}^{2D}_{y}\underline{u},$ (43)
$\underline{\textbf{curl}}^{y}\left(\underline{u}\right)$ $=-\underline{\Delta}^{2D}_{x}\underline{u},$ (44)
[…]
@@
$\underline{\text{curl}}\circ\underline{\textbf{grad}}\left(\underline{u}\right)$ $=\left(\underline{\Delta}^{2D}_{y}\underline{\Delta}^{2D}_{x}-\underline{\Delta}^{2D}_{x}\underline{\Delta}^{2D}_{y}\right)\underline{u}=\underline{0},$ (47)
$\underline{\text{div}}\circ\underline{\textbf{curl}}\left(\underline{u}\right)$ $=\left(\underline{\Delta}^{2D}_{x}\underline{\Delta}^{2D}_{y}-\underline{\Delta}^{2D}_{y}\underline{\Delta}^{2D}_{x}\right)\underline{u}=\underline{0},$ (48)
@@
With the matrix representation of our derivative operators, we have the commuting 2D de Rham complex of Fig. 5, using reduction operators $R^{0},\textbf{R}^{1}$ and $R^{2}$. These operators are defined as the mappings sending functions in the discrete spaces to the vector of their degrees of freedom. We can again evaluate the operators nodally using the histopolation Vandermonde matrix $\underline{V}$ analogously to the 3D case in equations (25) to (34).
@@
Figure 5: Discrete de Rham complex in 2D
$H^{1}$$H\left(\text{curl}\right)$$H\left(\text{div}\right)$$L^{2}$gradcurldiv
Figure 6: 3D de Rham complex
We can further prove that the cohomology spaces in the discrete setting are the same as in the continuous setting for non-periodic rectangular domains:
Proposition 1.
For a non-periodic rectangular domain all cohomology spaces of the discrete versions of the de Rham complexes without imposed boundary conditions in Fig. 2 and 3 vanish:
[…]
where $I\left(\mathbb{R}\right)$ is the inclusion of $\mathbb{R}$ into $W^{0}$ by mapping $c\in\mathbb{R}$ to the function with all coefficients equal to $c$, and $\tilde{W}^{2}$ is the space discretizing H(curl). In the 3D case, we have for the discrete versions of the 3D de Rham complex in Fig. 6:
[…]
where $W^{3}$ is the space discretizing $L^{2}$.
Proof:
The proof works by counting degrees of freedom. For details see appendix A.
@@ 2.3 Summation-By-Parts Finite Difference Operators
In the previous section, we assumed the existence of Lagrange functions that fulfill (2) and (3) at grid points $\{x_{i}\}_{i=0}^{N}$ (including the boundary nodes), but we did not discuss how to construct them. In [27], Gerritsma uses the Legendre–Gauss–Lobatto points and the Lagrange interpolating polynomials for those points. The corresponding derivative matrix $\underline{D}$ together with the mass matrix $\underline{M}$, which is a diagonal matrix with the Legendre–Gauss–Lobatto quadrature weights on the diagonal, has the summation-by-parts (SBP) property,
@@
As we have pointed out above, we only need to evaluate our basis functions and their derivatives at the grid nodes to construct the mimetic operators defined in the previous section. If we use a collocation quadrature rule when applying Finite Element Exterior Calculus [2, 3] to derive our scheme, the global shape of the basis functions $l_{i}$ does not matter at all, only their nodal function and derivative values are relevant. The central idea of this paper is therefore to not construct basis functions $l_{i}$ at all, but use existing diagonal norm finite difference SBP operators, for example from Strand [44], to supply a quadrature rule and nodal derivative data needed for the numerical scheme. We do this explicitly by prescribing the nodal derivatives of the Lagrange functions, that are not known globally, through the derivative matrix according to (13) and using diagonal entries of the norm matrix $\underline{M}$ as weights for a collocation quadrature rule on the grid points.
One of the standard SBP finite difference operators from [44], which we will use in this study, is the operator with interior order $4$ and boundary order $2$ with $N+1$ points:
@@
which is defined on a regular grid, where $\Delta x$ is the distance between two neighboring grid nodes. In the derivative matrix, the upper left boundary block is repeated in the bottom right with reversed signs and inverted order.
@@ 3.1 Maxwell’s Equations in 2D
Our goal is to derive a mimetic scheme using formulations analogous to finite element exterior calculus to preserve the divergence involution (66). For simplicity we will consider a rectangular domain, using a Cartesian mesh for the discretization. To apply the framework of Finite Element Exterior Calculus, we need the de Rham complexes containing the derivative operators of the PDEs.
@@ 3.2 De Rham Complexes
$H_{0}^{1}$$H_{0}\left(\text{div}\right)$$L^{2}$curldiv
Figure 7: 2D de Rham complex with homogeneous boundary conditions
$L^{2}$$H_{0}\left(\text{curl}\right)$$H_{0}^{1}$curl-grad
Figure 8: 2D dual de Rham complex with homogeneous boundary conditions
Maxwell’s equations have an inherent geometric structure described by the de Rham cochain complex of Fig. 2 and its dual in the periodic case in Fig. 3 in 2D. We define the spaces needed for the complexes by
@@
The complexes are dual for periodic domains in the sense that
@@
$\left<\text{div}\left(\boldsymbol{\Psi}\right),\phi\right$> $=\left<\boldsymbol{\Psi},-\textbf{grad}\left(\phi\right)\right$>, $\qquad\forall\boldsymbol{\Psi}\in H\left(\text{div}\right),\phi\in H^{1},$ (70)
where the brackets denote the $L^{2}$ inner product. In the non-periodic case, the duality equations (69) and (70) do not hold any longer due to appearing boundary integrals when integrating by parts. However, we can restore duality when considering homogeneous boundary conditions for the primal or the dual complex. We consider the spaces
$H^{1}_{0}$ $=\{w\in H^{1}\hskip 1.42271pt|\hskip 1.42271ptu_{|\partial\Omega}\equiv 0\},$ (71)
$H_{0}\left(\text{curl}\right)$ $=\{\textbf{w}\in H\left(\text{curl}\right)\hskip 1.42271pt|\hskip 1.42271pt\textbf{w}_{|\partial\Omega}\cdot\tau\equiv 0\text{ for all tangent vector fields }\tau\},$ (72)
$H_{0}\left(\text{div}\right)$ $=\{\textbf{w}\in H\left(\text{div}\right)\hskip 1.42271pt|\hskip 1.42271pt\textbf{w}_{|\partial\Omega}\cdot\textbf{n}\equiv 0\text{ for all normal vector fields }\textbf{n}\},$ (73)
where we again omit the region $\Omega$ in the space names for brevity. Note that there is no adjustment for the $L^{2}$ space, since the trace operator is not well-defined on this space. Imposing homogeneous boundary conditions on either the primal or the dual complex restores duality. In other words, the primal complex without imposed boundary conditions in Fig. 2 and the dual complex with homogeneous boundary conditions in Fig. 8 are dual to each other. So are complexes in Fig. 7 and in Fig. 3. If we now demand in equations (69) and (70) that $\psi\in H^{1}_{0}$ and $\boldsymbol{\Psi}\in H_{0}\left(\text{div}\right)$ or $\boldsymbol{\Phi}\in H_{0}\left(\text{curl}\right)$ and $\phi\in H^{1}_{0}$, the equations hold for non-periodic domains. The homogeneous boundary conditions for the primal complex are:
$B^{z}_{|\partial\Omega}$ $\equiv 0,$ (74)
$\textbf{E}_{|\partial\Omega}\cdot\textbf{n}$ $\equiv 0,$ (75)
for all normal vector fields n, while for the dual complex they are
$\textbf{E}_{|\partial\Omega}\cdot\textbf{v}\equiv 0,$ (76)
for all tangent vector fields v. We can also impose homogeneous boundary conditions on the primal complex for one part of the boundary and on the dual complex on the remaining part. We have to choose in which spaces to look for E and $B^{z}$, since we have to choose which complex to discretize. If we choose $\textbf{E}\in H(\text{div})$ or $\textbf{E}\in H_{0}(\text{div})$ and $B^{z}\in H^{1}$ or $B^{z}\in H_{0}^{1}$, we can solve Ampère’s law and Gauss’s law strongly, while we have to use Stokes’ theorem to solve Faraday’s law weakly. Analogously, if we choose $\textbf{E}\in H(\text{curl})$ or $\textbf{E}\in H_{0}(\text{curl})$ and $B^{z}\in L^{2}$, we can solve Faraday’s law strongly, but need to solve Ampère’s law and Gauss’s law weakly.
Since the divergence involution for the magnetic field falls away in the 2D transverse electric case and we want to preserve Gauss’ law for the electric field (66), we will choose $\textbf{E}\in H(\text{div})$ and $B^{z}\in H^{1}$, leading to the formulation for the periodic case:
@@
$\left<\frac{\partial{\textbf{E}}}{\partial{t}},\boldsymbol{\Phi}\right>-\left<\textbf{curl}\left(B^{z}\right),\boldsymbol{\Phi}\right$> $=0,\qquad$ $\forall\boldsymbol{\Phi}\in H(\text{div}),$ (78)
@@
For non-periodic problems, the function spaces must be modified according to the imposed boundary conditions. Boundary conditions imposed directly through the definition of the trial and test spaces correspond to essential boundary conditions. In contrast, boundary conditions arising through the boundary terms generated by integration by parts correspond to natural boundary conditions.
The finite element exterior calculus ansatz is to choose a discrete subcomplex of the de Rham complex, meaning we choose $W^{0}\subset H^{1}$, $W^{1}\subset H\left(\text{curl}\right)$, $W^{2}\subset L^{2}$ such that $\textbf{curl}\left(W^{0}\right)\subset W^{1}$ and $\text{div}\left(W^{1}\right)\subset W^{2}$. If we use essential boundary conditions, we instead choose subspaces of the spaces of the complex with homogeneous boundary conditions. In addition we need bounded projections $p_{0},\textbf{p}_{1}$ and $p_{2}$, so that the diagram of Fig. 5 commutes. We then restrict both our ansatz and our test spaces to the corresponding finite dimensional subspaces. This means we again have two ansatz spaces for the electric and magnetic field: We look for a numerical solution $\textbf{E}\in[0,T]\times W^{1}$ and $B^{z}\in[0,T]\times W^{0}$ so that
$\left<\frac{\partial{B^{z}}}{\partial{t}},\psi\right>+\left<\textbf{E},\textbf{curl}\left(\psi\right)\right$> $=0,\qquad$ $\forall\psi\in W^{0},$ (80)
$\left<\frac{\partial{\textbf{E}}}{\partial{t}},\boldsymbol{\Phi}\right>-\left<\textbf{curl}\left(B^{z}\right),\boldsymbol{\Phi}\right$> $=0,\qquad$ $\forall\boldsymbol{\Phi}\in W^{1},$ (81)
$\left<\text{div}\left(\textbf{E}\right),\mu\right$> $=0,\qquad$ $\forall\mu\in W^{2}.$ (82)
Since $\text{div}(\textbf{E})\in W^{2}$ we do get $\text{div}(\textbf{E})\equiv 0$ from (82). We will later show that by solving (80) and (81) with this method, we automatically preserve the divergence involution both in the semi-discrete and fully discrete setting. Therefore we will only consider the evolution equations (80) and (81) when deriving our method.